35. Demand for a managerial economics text is given by Q = 20,000 – 300P. The book is initially priced at $30:

i) Compute the point price elasticity of demand at P= $30.

ii) If the objective is to increase total revenue, should the price be increased or decreased? Explain.

iii) Compute the arc price elasticity for a price decrease from $30 to $20.

iv) Compute the arc price elasticity for a price decrease from $20 to $15.
1
Expert's answer
2020-12-02T09:59:29-0500

i) The point price elasticity of demand at P= $30 is: "Ed = -300\u00d730\/11,000 = -0.82."

ii) If the objective is to increase total revenue, the price should be increased, because the demand is inelastic.

iii) The arc price elasticity for a price decrease from $30 to $20 is:

"Ed =\frac{14,000-11,000} {20-30} ×\frac{20+30} {14,000+11,000} = -0.6."

iv) The arc price elasticity for a price decrease from $20 to $15 is:

"Ed =\frac{15,500-14,000} {15-20} ×\frac{15+20} {15,500+14,000} = -0.36."